OpenStudy (fredoneedshelp):

Translate into multi-step equation Twice the sum of a number and 5 equals 9 Use the variable c for the unknown number.

OpenStudy (sweetnikkie072):

can you revise your question..guess there is some error is it?

OpenStudy (fredoneedshelp):

Thats how it is

OpenStudy (ellalaynemaguire):

first of all, the number plus 5 is -0.5 second, the equation would be 2(5+c)=9

OpenStudy (fredoneedshelp):

Thanks man

OpenStudy (fredoneedshelp):

im falling behind on my work

OpenStudy (ellalaynemaguire):

me too

OpenStudy (ellalaynemaguire):

and no problem

OpenStudy (fredoneedshelp):

i need 3 right to go on the next one

OpenStudy (fredoneedshelp):

they're all practice for translating multi step

OpenStudy (ellalaynemaguire):

i can help with more if you need it, those kind of problems are really easy for me

OpenStudy (fredoneedshelp):

okay yeah, much appreciated bro

OpenStudy (fredoneedshelp):

The difference of a number divided by 8 and 4 is 9. Use the variable x for the unknown number.

OpenStudy (ellalaynemaguire):

(x÷8)-(x÷4)=9

OpenStudy (ellalaynemaguire):

is that what you needed?

OpenStudy (fredoneedshelp):

it doesnt show the division sign @ellalaynemaguire

OpenStudy (ellalaynemaguire):

okay well then the / will represent a division sign (x/8)-(x/4)=9